OpenStudy (anonymous):

Evaluate 8 x minus eleven for x equals sixteen. nineteen forty eighty-eight one hundred seventeen

OpenStudy (accessdenied):

In mathematical symbols, "8x-11", for x=16 Well, we replace x with 16 and evaluate. 8(16)-11 = 128 + 11 = 117 One-hundred Seventeen, the last answer.

OpenStudy (anonymous):

thankz

OpenStudy (accessdenied):

the wrong sign is used in the second line (128+11), should be 128-11. Sorry. :p

OpenStudy (anonymous):

so the last one is rong

OpenStudy (anonymous):

The answer is right, but AccessDenied was just saying that he wrote the problem wrong when he was showing his steps. 117 is still right.